In an 8 inch square cake pan and a 10 inch square cake pan, what is the difference in volume each pan will hold? Assume each pan is 1.5 inches high. Show step by step work. Include correct units with your solution.
Answer by edjones(7569)

V=s^2h
8^2*1.5=96
10^2*1.5=150
150-96=54 in^3
.
Ed